2025-08-29 On-sky Commissioning Update

During weeks 18-19 of the on-sky commissioning campaign with LSSTCam the team persisted with on-sky and in-dome engineering tests to improve the consistency of delivered image quality. Winter weather continued to impact the campaign, allowing only 4 nights of on-sky testing from 15-27 August, with more snow and ice, and temperatures at Cerro Pachón falling to as low as -8 C.

  • Active Optics System (AOS) commissioning included a series of tests to optimize performance of the closed loop control system:
    • Wavefront repeatability tests using both the corner wavefront sensors and while pistoning the full science array to characterize statistical noise in the wavefront measurements, sampling a range of conditions
    • Offline analysis and on-sky testing considering combinations of degrees of freedom to control specific wavefront aberrations expressed in a basis of Zernike polynomials (corresponding to, e.g., astigmatism, coma, trefoil, spherical)
    • Testing a new closed-loop configuration with more degrees-of-freedom enabled
  • Work on the primary-tertiary mirror (M1M3) thermal control systems included in-dome tests to calibrate the >150 temperature sensors for the M1M3 glass and testing of the thermal response when running the Fan Coil Units with a range of fan speeds. The team also conducted a first set of on-sky tests running the AOS closed loop at fixed telescope position while adjusting the thermal setpoint for M1M3 to lower and raise temperature relative to ambient. There are ongoing studies to correlate temperatures within the mirror, and differences with respect to dome ambient temperatures, with the delivered image quality, considering effects of both mirror seeing and thermally induced distortions to the M1M3 optical figure.
  • Daytime engineering work included preparations to install and use a first set of louvers to improve dome ventilation, as well as preparations to install sonic anemometers to monitor the dome environment.
  • There was limited progress on the SV surveys during the past two weeks, with ~650 science program visits acquired from 15-27 August. Rubin Observatory will continue to post updates for the SV survey progress and provide Nightly observing reports. Considering reduced opportunities for SV survey observations over the past month, Rubin Observatory is considering to prioritize a subset of the Deep Drilling Fields during the next weeks to increase the likelihood of achieving deep integrated exposure approaching the LSST 10-year equivalent depth to facilitate commissioning of the LSST Science Pipelines
  • In-dome testing used the collimated bream projector (CBP) to measure crosstalk coefficients and to re-produce some stay light features
  • Substantial progress was made on understanding the correlation between ambient temperature and the LSSTCam cryogenic cooling capacity, in particular, identifying condensation in the vapor discharge line as the source of the drop in cooling capacity. By reducing the temperature of the glycol which cools the phase separator in the cryo cabinet, the team was able to demonstrate stable operation of LSSTCam at lower ambient temperatures than previously achieved. Work continues on applying heat and insulation to the cryo refrigeration transport tubes as additional mitigation measures.
  • As noted in the previous update, the currently available filter set is griz, with one filter socket currently non-operations. No filter swaps are planned during the next two weeks.

There are approximately three weeks of the on-sky commissioning campaign remaining before an engineering and maintenance period currently scheduled for mid-September through mid-October.